### v1

@@ 3.3 Morphology and radial profile of the halo-like excess
Figure 9 shows the radial angular profiles of three halo model fits, at the energy bin of 21 GeV where the halo component is the most prominent. The diffuse flux of the best-fit halo model is shown as a curve, and the data points are the sum of the best-fit model and the fit residuals, in a ring region centered on the GC within the ROI (, ). This plot confirms what we found in the previous section, namely that the NFW- model best fits the data in terms of likelihood. The density profile of the NFW- model is steeper than the data, while the NFW- model is shallower. For comparison, the best fit flux of the GC GeV excess (the NFW- model in the fit including the Galactic plane) is also shown in the figure as the curve of “NFW- ()”. The halo component must have a shallow profile that does not exceed this curve when extrapolated to the GC. Though the NFW- model best fits at , it is several times higher than the GC excess at around 1∘ from the GC. This suggests that the angular profile that best fits the data across the entire ROI is slightly flatter than NFW-, or that the angular profile becomes gradually flatter than NFW- toward the GC.
@@
To verify how well the halo model matches the data in terms of morphology, map images are shown in figure 10, for the NFW- fit at the 21 GeV bin. When the fit is made without a halo model component, the fit tries to minimize the halo component, and hence, the fit would be biased if the halo component physically exists. However, a halo-like residual can be seen even in such a fit (top left). In the fit including NFW- as the halo model, the map of the model plus residual (i.e., the observed map with all model maps except for the halo subtracted) shows a clearer halo-like structure (top right), which is spherically symmetric and consistent with the model map (bottom left). The fit residual with the halo model (bottom right) is significantly reduced compared to the fit without the halo (top left), indicating that the halo fit is successful overall in the ROI.
@@
To examine the fits in energy bins other than the peak (21 GeV), figures 11 and 12 show maps of the NFW- model plus residual at bins lower and higher than 21 GeV, respectively. These maps show the trend seen in figure 7, with the negative fit at the lowest 1.5 GeV, nearly zero at 2.5 GeV, the halo component becoming stronger toward 21 GeV, and then rapidly weaker beyond 21 GeV. The morphology of the halo model plus residual is generally spherically symmetric and does not appear to contradict the profile of the assumed model halo.
@@ 3.4 Examination of systematic uncertainties
Here, we examine systematic uncertainties concerning the evidence for the halo-like excess on the following points: (1) uncertainties in the Fermi bubble template, (2) treatment of point sources, (3) model parameters and number of free components in GALPROP, and (4) halo-like excess search using the Fermi LAT standard background model (GIEM) as the reference.
@@ 3.4.1 The Fermi bubble template
The Fermi bubble template was created from the positive residual map at 4.3 GeV; however, a portion of the halo component in that energy bin may be absorbed by the template, resulting in an underestimate of the halo component in the fit. To check the dependence on the choice of energy bin in this process, an alternative fit result is shown in Figure 14 (top), in which the positive (the Fermi bubbles) and negative residual templates are replaced with those of the 1.5 GeV map shown in figure 1. As expected, the energy bin where the halo component becomes zero has shifted to the lower side, and the halo flux at the 20 GeV peak has increased, but there is no change in the basic shape of the spectrum with a peak near 20 GeV.
As shown in figure 1, due to the strong negative residual around 10–15∘, the positive residual template may be biased from the true structure of the Fermi bubbles. To test the dependence on the Fermi bubble template structure, the fit result obtained by replacing the positive residual template with the simple flat template is shown in Figure 14 (top). The halo component has decreased compared to the baseline analysis, and this may be because the baseline analysis attempted to compensate for the concavity caused by the negative residual with the halo component. However, the overall spectral shape, including the rising rate of the halo component from 1 to 20 GeV, remains unchanged. The excess at the 21 GeV bin is still statistically significant (9.1).
@@ 3.4.2 Point sources
In the baseline analysis, point sources are subtracted using the spectral models in the 4FGL-DR4 catalog and the LAT PSF. Instead, the fit result after masking circular areas around the point sources with a radius equal to the 95% containment angle is shown in figure 14 (top). The result is almost unchanged from the baseline analysis. Due to the deterioration of the LAT angular resolution, the area to be masked increases toward lower photon energies, resulting in a large statistical error at the lowest bin.
@@ 3.4.3 The GALPROP models
In the LAT team paper on the GC GeV excess [7], the GALPROP gas component is divided into five regions of Galactocentric radius (0–1.5, 1.5–3.5, 3.5–8, 8–10, and 10–50 kpc), and the ICS component is divided into the three sub-components of the different target photon radiation fields (optical, infrared, and CMB), which are fitted as independent templates. To check the dependence on the number of free GALPROP components, the NFW- fit is performed with the same treatment about GALPROP as ref. [7], and the result is shown in Figure 14 (top). The halo component decreases in bins below 12 GeV from the baseline analysis, but remains almost unchanged at the 21 GeV bin, and the overall trend of the spectrum peaking at 20 GeV remains unchanged.
@@ 3.4.4 The GIEM as the zero point
The reason is that structures such as the Fermi bubbles, the GC excess, and Loop I, for which it is difficult to create reliable templates from other wavelength data, are modeled by iteratively generating non-template patches from the fit residuals.
@@
This test is performed by a fit including the GIEM and isotropic components with non-zero initial values in MCMC (the original model normalization for the GIEM), and the NFW- model with zero initial value. However, with this method, any map component that cannot be described by a single power law spectrum will remain as an excess. To avoid confusion between the halo and other components, the other model components used in the baseline analysis (GALPROP gas and ICS, Loop I, and the structured Fermi bubble template) are also included in the fit with zero initial values in MCMC. The halo component spectrum obtained by this fit is shown in figure 14 (top). The halo component decreases compared to the baseline analysis, but is still statistically significant with the same peak at around 20 GeV (6.7, 7.4, and 5.1 at 12, 21, and 35 GeV bins, respectively). Since a part of the halo component may have been reduced by the non-template patch in the GIEM, this result can be considered as the lower limit for the halo component.
@@ 4.2 Fitting of the WIMP parameters
Fits for the four cases (three for NFW-, and one for NFW-) are shown in figure 15, where the best-fit WIMP parameters are also shown. As photon energy decreases, statistical errors in the data become smaller, and hence spectral fitting tends to be influenced by the lowest energy bins. However, as we have already seen, the systematic error is much larger than the statistical error in these bins. Therefore, the WIMP parameters obtained here should be considered as rough estimates based on the peak position and normalization of the halo excess. For this reason, the statistical errors of the WIMP parameters are not reported. The NFW- halo flux of the baseline analysis is negative in the lowest energy bin, and the WIMP parameter fit is strongly affected by this negative value and its small statistical error. Therefore, from among the alternative analyses relative to the NFW- baseline, three are selected in which the halo flux of the lowest energy bin is not strongly negative: (1) the SNR distribution for cosmic-ray sources in GALPROP, (2) the Fermi bubble template created at 1.5 GeV, and (3) the GIEM used as the reference in the halo excess search. The case (2) results in the highest halo component flux (and therefore the largest cross section), while the case (3) results in the lowest. For the reason stated in section 3.4.4, the cross section of the case (3) should be considered as a lower limit.
In the NFW- fits, 0.4–0.8 TeV and (5–9) are favored for the and channels, and 40–80 GeV and (7–9) for the channel. The cross sections favored in the NFW- fit are ( and ) and (). These correspond to the decay timescale and s, respectively, when dark matter decays rather than annihilates. In the fits for the channel, is likely biased toward a lower value because of the small statistical errors in low-energy bins. The halo excess spectrum seems to fit better the and channels than due to their broader annihilation spectra, though the halo excess spectrum contains significant systematic uncertainties, particularly in the low-energy region.
@@ 4.3.1 Relationship with the GC GeV excess
It should be noted that the halo excess has a flux similar to that of the GC excess at and around 2 GeV, where the GC excess peaks. If past studies have confused part of the halo excess with the extended tail of the GC excess, the true GC excess may be more centrally concentrated than . Extension of the GC excess to more than 10∘ has been argued to support the dark matter origin [62, 63, 64], but if the true GC excess is more centrally concentrated, it may be more consistent with stellar population origins, such as millisecond pulsars.
@@ 4.3.2 Comparison with previous constraints and theoretical expectations
The annihilation cross section estimated from the 20 GeV halo excess ( with GeV for the channel) is marginally consistent with the upper limits obtained by earlier studies on the MW halo region [26, 27, 28, 29, 31]. The LAT team’s analysis using two years of the LAT data [26] reported without background modeling, for their standard choice of the dark matter density at the solar neighborhood . It should also be noted that the ROI of the LAT team’s analysis (, ) is much narrower in the Galactic latitude than that of this study.
The cross section required for the halo excess is larger than the upper limits by dwarf spheroidal galaxies, which are currently considered to be the strongest [10, 11, 12, 13]. For example, it is several times larger than the 95% containment region of the constraints obtained in ref. [13]. However, this tension does not immediately rule out the dark matter interpretation of the 20 GeV halo excess. Compared to dwarf galaxies, whose dark matter mass can be more reliably constrained by observations, the density distribution of the MW halo is highly uncertain ([4], see also section 4.3.3 below). The dark matter content in dwarf galaxies is also still subject to uncertainties regarding methodology and observed data [65, 66]. In fact, the dark matter interpretation of the GC GeV excess is also in the same degree of tension with dwarf galaxy constraints, but for the same reason, the dark matter interpretation is not considered to be excluded [8, 13].
@@ 4.3.3 Uncertainties in the MW dark matter density distribution
The uncertainties in the density profile of the MW halo can be described in more detail as follows. Since the mass of baryons exceeds that of dark matter in the inner region of the MW ( 10 kpc) [67, 68], we can only extrapolate from the dark matter density at the solar system () assuming a halo density profile, but is uncertain by a factor of a few. There are several other analytic functions proposed for the halo density profile, in addition to NFW, and the Einasto profile [69], which is known to fit numerical simulations as well as NFW, is particularly interesting for the halo excess. When is fixed, this profile predicts a higher density than NFW at around 10∘ ( kpc) from the GC, and a shallower profile than NFW near the GC (see figure 2.3 of [4]), which are in agreement with the trends suggested by the 20 GeV halo excess.
In addition to the uncertainty in the fitting formulae for the mean halo density profile in numerical simulations, individual halo profiles also vary depending on their formation and merger history. Furthermore, in the central region of a halo, the dark matter density profile may be affected by complicated baryon dynamics, and this effect is not yet fully understood [70, 71]. In the case of a smooth density distribution (), a threefold uncertainty in the estimate of dark matter density in the 10∘–20∘ region from the GC changes the halo flux by nearly an order of magnitude. Finally, the effect of subhalos or any substructures in the MW halo should not be overlooked. Since substructures are more easily destroyed in the central region [52], the signal may be boosted with an effective density profile shallower than NFW, in a direction consistent with the trends of the halo excess.
@@ 4.3.4 The extragalactic gamma-ray background radiation
If the MW halo emits annihilation gamma-rays, the extragalactic gamma-ray background radiation (EGB) should include a contribution from annihilations in all dark halos in the universe. Theoretical predictions suggest that the EGB flux by annihilation is comparable to that from the MW halo, with the peak photon energy redshifted to about half [22]. The observed EGB in the 1–1000 GeV range is thought to contain various astronomical components, such as active galactic nuclei and star-forming galaxies, making it difficult to identify the dark matter component unambiguously. A recent study suggests that a significant contribution to the EGB from annihilation with the WIMP parameters of the 20 GeV halo excess is not ruled out [25].
The EGB from star-forming galaxies is well understood in terms of physical processes, and relatively reliable theoretical models have been developed [23, 72]. These models contribute significantly to the Fermi LAT data below 10 GeV, but their predicted flux decreases rapidly above 10 GeV due to intergalactic absorption. On the other hand, the EGB data extend approximately as a power law beyond 10 GeV, and a new component may be appearing in this region. Annihilation gamma rays from the MW halo and extragalactic halos, whose spectral peak is around 10–20 GeV, may be a candidate for such a component. In fact, this was the original motivation behind this study to search for signs of dark matter annihilation.

### v2

@@ 3.3 Morphology and radial profile of the halo-like excess
Figure 10 shows the radial angular profiles of three halo model fits at the energy bin of 21 GeV, where the halo component is the most prominent. The diffuse flux of the best-fit halo model is shown as a curve, and the data points are the sum of the best-fit model and the fit residuals, in a ring region centered on the GC within the ROI (, ). It is not straightforward to determine from this figure which of the three halo models fits best, but as seen in the previous section 3.2 and figure 9, the likelihood values indicate that the NFW- model is preferred with a statistical significance of 2 or greater. The density profile of the NFW- model is steeper than the data, while the NFW- model is shallower.
The data points differ significantly among the three models, especially in the tail region, suggesting the presence of degeneracy with other model components. In particular, the tail of the halo models is considered prone to degeneracy with the isotropic component. Indeed, in the NFW- fit, the isotropic component decreases significantly near 20 GeV compared to the NFW- fit, becoming consistent with zero (figures 6 and 7). In the NFW- fit, the flux of the isotropic component near 20 GeV (figure 6) agrees well with the unresolved isotropic background flux reported by the Fermi-LAT team [56], (shown in figure 10). This implies that the NFW- fit leaves no room for the isotropic background that should inherently exist, providing further evidence that the NFW- model fits the data better, in addition to the likelihood values.
For comparison, the best fit flux of the GC GeV excess (the NFW- model in the fit including the Galactic plane, figure 2) is also shown in the figure as the curve of “NFW- ()”. The halo component must have a shallow profile that does not exceed this curve when extrapolated to the GC. Though the NFW- model best fits at , it is several times higher than the GC excess at around 1∘ from the GC. This suggests that the angular profile that best fits the data across the entire ROI is slightly flatter than NFW-, or that the angular profile becomes gradually flatter than NFW- toward the GC.
@@
To verify how well the halo model matches the data in terms of morphology, map images are shown in figure 11, for the NFW- fit at the 21 GeV bin. When the fit is made without a halo model component, the fit tries to minimize the halo component, and hence, the fit would be biased if the halo component physically exists. However, a halo-like residual can be seen even in such a fit (top left), giving a lower bound for the halo-like signal. In the fit including NFW- as the halo model, the map of the model plus residual (i.e., the observed map with all model maps except for the halo subtracted) shows a clearer halo-like structure (top right), which is spherically symmetric and consistent with the model map (bottom left). The fit residual with the halo model (bottom right) is significantly reduced compared to the fit without the halo (top left), indicating that the halo fit is successful overall in the ROI.
@@
To examine the fits in energy bins other than the peak (21 GeV), figures 12 and 13 show maps of the NFW- model plus residual at bins lower and higher than 21 GeV, respectively. These maps show the trend seen in figure 8, with the negative fit at the lowest 1.5 GeV, nearly zero at 2.5 GeV, the halo component becoming stronger toward 21 GeV, and then rapidly weaker beyond 21 GeV. The morphologies of the halo model plus residual are generally spherically symmetric and do not appear to contradict the profile of the assumed model halo.
@@ 3.4 Examination of systematic uncertainties
Here, systematic uncertainties concerning the evidence for the halo-like excess are examined on the following points: (1) fit analysis divided into four quadrants, (2) uncertainties in the Fermi bubble template, (3) treatment of point sources, (4) model parameters and number of free components in GALPROP, and (5) halo-like excess search using the Fermi LAT standard background model (GIEM) as the reference.
3.4.1 Analysis divided into four quadrants of the Galactic coordinates
To test the isotropy of the halo excess around the GC, the NFW- fits are performed by dividing into four quadrants about the Galactic coordinates, and the halo excess spectra are shown in figure 15 (top panel). Furthermore, the fitting results are also illustrated when dividing the four quadrants shifted by 45∘ into two quadrants in the galactic plane direction () and two quadrants in the polar direction (). Although statistical fluctuations increase significantly above 30 GeV due to the decrease in photon number, a consistent trend is observed in all quadrants below 20 GeV.
@@ 3.4.2 The Fermi bubble template
The Fermi bubble template was created from the positive residual map at 4.3 GeV; however, a portion of the halo component in that energy bin may be absorbed by the template, resulting in an underestimate of the halo component in the fit. To check the dependence on the choice of energy bin in this process, an alternative fit result is shown in figure 15 (middle), in which the positive (the Fermi bubbles) and negative residual templates are replaced with those of the 1.5 GeV map shown in figure 1. As expected, the energy bin where the halo component becomes zero has shifted to the lower side, and the halo flux at the 20 GeV peak has increased, but there is no change in the basic shape of the spectrum with a peak near 20 GeV.
As shown in figure 1, due to the strong negative residual around 10–15∘, the positive residual template may be biased from the true structure of the Fermi bubbles. To test the dependence on the Fermi bubble template structure, the fit result obtained by replacing the positive residual template with the simple flat template is shown in figure 15 (middle). The halo component has decreased compared to the baseline analysis, likely because the bubble flux in the negative residual region becomes larger compared to the structured bubble template, and the halo component is fitted smaller to compensate for this. However, the overall spectral shape, including the rising rate of the halo component from 1 to 20 GeV, remains unchanged. The excess at the 21 GeV bin is still statistically significant (9.1).
As a further verification of systematic uncertainties related to the negative residual region, the fit is performed after this region is masked. Pixels with notably negative flux () in the right panel of figure 1 are excluded from the ROI, and the negative residual map is removed from the set of the fitting templates. The fit result is shown in figure 15 (middle); the halo excess spectrum does not differ significantly from the baseline analysis.
@@ 3.4.3 Point sources
In the baseline analysis, point sources are subtracted using the spectral models in the 4FGL-DR4 catalog and the LAT PSF. Instead, the fit result after masking circular areas around the point sources with a radius equal to the 95% containment angle is shown in figure 15 (middle). The result is almost unchanged from the baseline analysis. Due to the deterioration of the LAT angular resolution, the area to be masked increases toward lower photon energies, resulting in a large statistical error at the lowest bin.
@@ 3.4.4 The GALPROP models
In the LAT team paper on the GC GeV excess [7], the GALPROP gas component is divided into five regions of Galactocentric radius (0–1.5, 1.5–3.5, 3.5–8, 8–10, and 10–50 kpc), and the ICS component is divided into the three sub-components of the different target photon radiation fields (optical, infrared, and CMB), which are fitted as independent templates. To check the dependence on the number of free GALPROP components, the NFW- fit is performed with the same treatment about GALPROP as ref. [7], and the result is shown in figure 15 (middle). The halo component decreases in bins below 12 GeV from the baseline analysis, but remains almost unchanged at the 21 GeV bin, and the overall trend of the spectrum peaking at 20 GeV remains unchanged.
@@ 3.4.5 The GIEM as the zero point
The reason is that diffuse structures such as the Fermi bubbles, the GC excess, and Loop I, for which it is difficult to create reliable templates from other wavelength data, are modeled by iteratively generating non-template patches from the fit residuals.
@@
This test is performed by a fit that includes the GIEM, point sources, and the isotropic background with non-zero initial values in MCMC (the original model normalization for the GIEM), as well as the NFW- model with a zero initial value. However, with this treatment, any map component that a simple power law spectrum cannot describe will remain as an excess. To avoid confusion between the halo and other components, the model components used in the baseline analysis (GALPROP gas and ICS, Loop I, and the structured Fermi bubble template) are also included in the fit with zero initial values in MCMC. The halo component spectrum obtained by this fit is shown in figure 15 (middle). The halo component decreases compared to the baseline analysis, but is still statistically significant with the same peak at around 20 GeV (6.7, 7.4, and 5.1 at 12, 21, and 35 GeV bins, respectively). Since a part of the halo component may have been reduced by the non-template patch in the GIEM, this result can be considered as a lower limit for the halo component.
@@ 4.2 Fitting of the WIMP parameters
Fits for the four cases (three for NFW-, and one for NFW-) are shown in figure 16, where the best-fit WIMP parameters are also shown. As photon energy decreases, statistical errors in the data become smaller, and hence spectral fitting tends to be influenced by the lowest energy bins. However, as we have already seen, the systematic error is much larger than the statistical error in these bins. Therefore, the WIMP parameters obtained here should be considered as rough estimates based on the peak position and normalization of the halo excess. For this reason, the statistical errors of the WIMP parameters are not reported. The NFW- halo flux of the baseline analysis is negative in the lowest energy bin, and the WIMP parameter fit is strongly affected by this unphysical negative value and its small statistical error. Therefore, from among the alternative analyses relative to the NFW- baseline, three are selected in which the halo flux of the lowest energy bin is not strongly negative: (1) the SNR distribution for cosmic-ray sources in GALPROP, (2) the Fermi bubble template created at 1.5 GeV, and (3) the GIEM used as the reference in the halo excess search. The case (2) results in the highest halo component flux (and therefore the largest cross section), while the case (3) results in the lowest. For the reason stated in section 3.4.5, the cross section of the case (3) should be considered as a lower limit.
In the NFW- fits, 0.4–0.8 TeV and (5–9) are favored for the and channels, and 40–80 GeV and (7–9) for the channel. The cross sections favored in the NFW- fit are ( and ) and (). These correspond to the decay timescale and s, respectively, when dark matter decays rather than annihilates. In the fits for the channel, is likely biased toward a lower value because of the small statistical errors in low-energy bins. The halo excess spectrum seems to fit the and channels better than due to their broader annihilation spectra, though the halo excess spectrum contains significant systematic uncertainties, particularly in the low-energy region.
@@ 4.3.1 Relationship with the GC GeV excess
It should be noted that the halo excess has a flux similar to that of the GC excess at and around 2 GeV, where the GC excess peaks (see figures 3 and 6). If past studies have confused part of the halo excess with the extended tail of the GC excess, the true GC excess may be more centrally concentrated than . Extension of the GC excess to more than 10∘ has been argued to support the dark matter origin [64, 65, 66], but if the true GC excess is more centrally concentrated, it may be more consistent with stellar population origins, such as millisecond pulsars.
@@ 4.3.2 Comparison with previous constraints and theoretical expectations
The annihilation cross section estimated from the 20 GeV halo excess ( with GeV for the channel) is marginally consistent with the upper limits obtained by earlier studies on the MW halo region [28, 29, 30, 31, 33]. The LAT team’s analysis using two years of the LAT data [28] reported without background modeling, for their standard choice of the dark matter density at the solar neighborhood, . It should also be noted that the ROI of the LAT team’s analysis (, ) is much narrower in the Galactic latitude than that of this study.
The cross section required for the halo excess is larger than the upper limits by dwarf spheroidal galaxies, which are currently considered to be the strongest [12, 13, 14, 15]. For example, it is several times larger than the 95% containment region of the constraints obtained in a recent study [15]. However, this tension does not immediately rule out the dark matter interpretation of the 20 GeV halo excess. Compared to dwarf galaxies, whose dark matter mass can be more reliably constrained by observations, the density distribution of the MW halo is highly uncertain ([4], see also section 4.3.3 below). The dark matter content in dwarf galaxies also remains subject to uncertainties regarding methodology and observed data [67, 68]. In fact, the dark matter interpretation of the GC GeV excess is also in a similar degree of tension with dwarf galaxy constraints, but for the same reason, the dark matter interpretation is not considered to be excluded [8, 15].
@@ 4.3.3 Uncertainties in the MW dark matter density distribution
The uncertainties in the density profile of the MW halo can be described in more detail as follows. Since the mass of baryons exceeds that of dark matter in the inner region of the MW ( 10 kpc) [69, 70, 71, 72], we can only extrapolate from the dark matter density at the solar system () assuming a halo density profile. In addition to the uncertainty of the density profile, the estimate for has an uncertainty of a factor of about two centered around [4]. In the case of a smooth density distribution (), a threefold uncertainty in the dark matter density at 10∘–20∘ from the GC changes the annihilation flux by nearly an order of magnitude.
There are several other analytic functions proposed for the halo density profile, in addition to NFW, and the Einasto profile [73], which is known to fit numerical simulations as well as NFW, is particularly interesting for the halo excess. When is fixed, this profile predicts a higher density than NFW at around 10∘ ( kpc) from the GC, and a shallower profile than NFW near the GC (see figure 2.3 of [4]), which are in agreement with the trends suggested by the 20 GeV halo excess. Some reports imply that a more core-like density profile than NFW better matches observational data for the MW halo [71, 72].
In addition to the uncertainty in the fitting formulae for the mean halo density profile in numerical simulations, individual halo profiles also vary depending on their formation and merger history. Furthermore, in the central region of a halo, dark matter density profile may be affected by complicated baryon dynamics, and this effect is not yet fully understood [74, 75]. Finally, the effect of subhalos or any substructures in the MW halo should not be overlooked. Since substructures are more easily destroyed in the central region [54], the signal may be boosted with an effective density profile shallower than NFW, in a direction consistent with the trends of the halo excess.
@@ 4.3.4 The extragalactic gamma-ray background radiation
If the MW halo emits annihilation gamma-rays, the extragalactic gamma-ray background radiation (EGB) should include a contribution from annihilations in all dark halos in the universe. Theoretical predictions suggest that the EGB flux by annihilation is comparable to that from the MW halo, with the peak photon energy redshifted to about half [24]. The observed EGB in the 1–1000 GeV range is thought to contain various astronomical components, such as active galactic nuclei and star-forming galaxies [25], making it difficult to identify the dark matter component unambiguously. A recent study suggests that a significant contribution to the EGB from annihilation with the WIMP parameters of the 20 GeV halo excess is not ruled out [27].
The EGB from star-forming galaxies is well understood in terms of physical processes, and relatively reliable theoretical models have been developed [25, 76]. These models contribute significantly to the Fermi LAT data below 10 GeV, but their predicted flux decreases rapidly above 10 GeV due to intergalactic absorption. On the other hand, the EGB data extend approximately as a power law beyond 10 GeV, and a new component may be appearing in this region. Annihilation gamma rays from the MW halo and extragalactic halos, whose spectral peak is around 10–20 GeV, may be a candidate for such a component. In particular, the emission from the MW halo is expected to be relatively stronger at higher photon energies because it is not subject to intergalactic absorption, and some of it may have been confused with the isotropic background radiation. In fact, this was the original motivation behind this study to search for signs of dark matter annihilation from the MW halo.